# Scattering and induced false vacuum decay in the two-dimensional quantum Ising model

### v3

@@ Abstract
Our results demonstrate that tensor networks can capture scattering dynamics and probe non-perturbative physics in 2D. This opens the door to such simulations in more complex settings, like lattice-gauge theories and condensed matter systems, and paves the way for future quantum simulations in regimes out of reach of classical simulations.
@@ I The model and its low-energy spectrum
The elementary excitations above the polarized ground state are magnons; spin flips generated by the $X$ operator. Their excitation energy is $8J$ for $g=h=0$, and they acquire a finite dispersion with at finite $g/J$. At leading order the dispersion relation reads [40]:
$\varepsilon_{1}=8J-\frac{g^{2}}{J}\left(\frac{1}{2}+\frac{1}{4}\left(\cos k_{x}+\cos k_{y}\right)\right)\,.$ (2)
@@ II From elastic to inelastic scattering
Figure 2: Scattering of magnon wave packets. (a) The evolution of energy density for varying $g/J$. The wave packets have momenta $\mathbf{k}=\pm\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (b) The evolution of energy density at $g/J=1.5$ for two values of $\mathbf{k}=\pm(k,k)$. (c) The evolution of the quasiparticle correlators for representative values of $g/J$. The value at $tJ=10$ is subtracted to disentangle the scattering from the background effects. (d) The $g$-dependence of the peak value of the correlators. (e) The long-range correlations. The points where the correlations are measured are sketched in the top right panel of (a) using the same colors scheme. As they are smaller than the two-point measurements, the three-point correlators are multiplied by a factor of 3 for visibility. (f) The $g$-dependence of the peak value of the long-range correlations. The curves drawn with different shades of the same color correspond to simulations with different bond dimension, the darkest being the largest. We use 100 and 150 in the left, 100, 150, 200 in the left central, and 100, 150, 200, 250 in the two panels on the right.
@@
To visualize the wave packets, we use the energy density $\varepsilon_{i}\coloneqq-\frac{1}{2}J\sum_{j}Z_{i}Z_{j}-gX_{i}-hZ_{i}$ with $j$ neighbors of $i$. Its evolution through the scattering process is shown in Fig. 2(a) for three representative values of $g/J$ and $\mathbf{k}=\pm\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. The scattering process is predominantly elastic for $g/J=1$ (top). At $g/J=1.5$ (middle) it goes through an intermediate resonance, and we find signatures of production of a heavier particle for $g/J=2.0$ (bottom).
@@
To quantify the particles produced during the scattering process, we measure correlators corresponding to approximate creation operators of composite excitations, built from products of the number operator $n_{i,j}=\frac{1}{2}\left(Z_{i,j}+\mathds{1}_{i,j}\right)$ on neighboring lattice sites. We take the sum across all lattice sites. The three-spin horizontal ($h$) and kink ($k$) excitations are measured by
$\begin{split}C_{3}^{h}&=2\sum_{ij}\langle(\mathds{1}-n)_{i-2,j}n_{i-1,j}n_{i,j}n_{i+1,j}(\mathds{1}-n)_{i+2,j}\rangle,\\ C_{3}^{k}&=4\sum_{ij}\langle n_{i,j}n_{i+1,j}n_{i,j+1}(\mathds{1}-n)_{i+1,j+1}\rangle.\end{split}$ (3)
@@
The integer prefactors of $C_{3}^{h}$ and $C_{3}^{k}$ account for the degeneracy of these excitations; the signal of $C_{3}^{h}$ matches with the one of vertically oriented three-spin excitations, and the same for the four differently oriented three-spin kinks. The $(\mathds{1}-n)$ terms are necessary to disentangle the signals from larger particles. Without them, $C_{3}^{k}$ would also partially capture the four-spin particles described by $C_{4}$.
At finite $g$ these correlators do not measure the quasiparticles exactly, but only their approximations. One way to amend this would be to adiabatically turn off the transverse field after the scattering [10], to reinstate the product-state nature of the excitations. In contrast with 1D cases, this approach does not seem feasible in our case. The timescale for an accurate reverse quench is much larger than the decay time of heavier products, and more importantly, the time needed for the products to reach the edge of the system. Hence, we disentangle the scattering signal from the background by subtracting the value of the correlators at the end of the preparation procedure, at $tJ=10$.
@@
The evolution of the correlations is shown in Fig. 2(e), and $g$-dependence of peak amplitudes in Fig. 2(f). At $g/J=1$, the single dominant peak is a signature of the elastic process. It splits into two with increasing $g$, and at $g/J\gtrsim 1.5$ we clearly find two separate contributions. The first still corresponds to the elastic process, while the second signals the delayed arrival of particles generated through an intermediate resonance. The second peak is accompanied by a peak in $c_{O}$. The peaks of $c_{A}$ and $c_{O}$ are not equal because the magnon dispersion is not exactly sinusoidal, and thus the scattering is not precisely at the van Hove singularity. Note that the products are more concentrated in the off-axis direction for $g/J=1.5$. We expect that this is accompanied by a finite-size effect, as the products are not point-like particles, but rather extended objects on the lattice. The negative $c_{AO}$ that accompanies the overlap of $c_{A}$ and $c_{O}$ shows that the presence of particles on the two axis is anti-correlated. This confirms that these are two separate two-particle processes, and not a four-particle process.
@@ III Scattering in a false vacuum
Figure 3: Scattering in the false vacuum. (a) A sketch of the energy landscape of a true vacuum bubble with radius $r$. (b) The evolution of the energy density during scattering with $h/J=0.1$ and $h/J=0.3$, both with $g/J=1.5$ and $\mathbf{k}=\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (c) The change in magnetization per site with varying $h$ for $g/J=1.5$ and $\mathbf{k}=\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (d) The dependence of the threshold value $h^{*}$ on $g/J$. The gray dashed line corresponds to the threshold $h^{*}$ expected for $g\rightarrow 0$. See text for details. (e) The evolution of the radius of the true vacuum bubble in log-log scale for varying $g$, with some longitudinal field larger than the threshold value. We subtract the time $t_{0}$ when the radius reaches $r_{0}=3$ on the horizontal, and $r_{0}$ on the vertical axis. The black dashed line corresponds to linear growth and acts as a guide to the eye. (Inset) The evolution of the radius in lin-lin scale. In (c) and (e), the curves with different shades of the same color correspond to simulations with different bond dimension, darker is bigger. Here we use 150 and 200.
@@
Originating in cosmology and high-energy physics [42], the question has recently attracted attention in the context of metastability in quantum many-body systems. It has been studied in 1D numerically [43, 44, 28, 45, 46] and experimentally [47, 48, 49, 50]. The microscopic mechanism behind the decay is probabilistic bubble nucleation [45]: the vacuum fluctuations induced through a quench are small bubbles of the true vacuum, where the system locally tunnels into the stable state. Assuming a spherical bubble with radius $r$, its energy consists of two contributions: a positive interface tension, and a negative bulk term coming from the energy difference between the two vacuua [42]. In 2D one expects the energy $E(r)=A\cdot 2J\cdot r-B\cdot 2h\cdot r^{2}$, with model-dependent geometrical factors $A$ and $B$. The parabola acts as a potential barrier between the false vacuum at $r=0$ and the true vacuum at $r\rightarrow\infty$, with the maximal energy at the critical bubble size $r^{*}\propto J/h$ and $E(r^{*})\propto J^{2}/h$. If a bubble bigger than $r^{*}$ is created, it can expand by dissipating energy, and the system transitions into the stable manifold. The energy manifold is sketched in Fig. 3(a).
@@
We aim to induce this process through collisions of magnons – true vacuum bubbles – on the false vacuum background. The evolution of the energy density is shown in Fig. 3(b). By varying the threshold energy $E(r^{*})$ through $h$ at a fixed $g$, we find two sharply separated dynamical regimes. At small $h$ we recover the process akin to the one in Fig. 2. However, if $h$ is tuned beyond a threshold value, denoted $h^{*}$, the dynamics drastically changes. We find that the scattering process generates a bubble of the true vacuum, which violently expands and encompasses the whole system. This is reflected in the evolution of the magnetization, shown in Fig. 3(c), exhibiting two classes of behavior depending on $h$: it is approximately constant at $h<h^{*}$, and rapidly increases after scattering when $h>h^{*}$.
At first sight, this behavior seems intuitive; if the wave packets carry enough energy to classically cross the potential barrier a critical bubble can be created. However, the result is not in agreement with the existing literature on induced false vacuum decay in quantum field theory. Whereas the false vacuum decay can be induced in various ways, it was shown that a collision of two particles should not trigger it: semi-classical calculations imply that the scattering cross-section for such process remains exponentially suppressed up to infinite energy [36, 37, 38]. A simple explanation is that while the scattering might carry enough energy to create the true vacuum bubble, it is a state of many excitations, and generating it from a two-particle state is an unlikely high-order process. We thus attribute our observation of an induced false vacuum decay to the fact that we are working with larger values of $g/J$, possibly outside of the validity of the semi-classical arguments of Refs. [36, 37, 38], and where non-perturbative coherent processes become available. This conclusion is supported by the dependence of the threshold $h^{*}$ on $g$, shown in Fig. 3(d). The dashed line corresponds to the value of $h^{*}\sim 0.39$ expected in the classical point, $g\rightarrow 0$, obtained by equating the scattering energy $2\times 8J$ with $E(r=r^{*},h^{*})$. We assume a spherical bubble in 2D, hence $A=2\pi$, $B=\pi$. If the induced decay were only conditioned on the energy available, the curve would tend towards this value at small $g$. However, we find that as $g$ decreases $h^{*}$ rapidly increases. This indicates that the decay is induced only for a small potential barrier (small $E^{*}$), where the system can tunnel through.
@@
Finally, in Fig. 3(e), we plot the evolution of the bubble radius for various $g/J$, with $h>h^{*}(g)$. The main panel shows a log-log plot where we subtract a $g$-dependent $t_{0}$ from the time axis, and $r_{0}=3$ in the vertical axis. The inset shows the raw data in linear scale. The black dashed line corresponds to linear increase of $r$. It roughly matches the numerical data, implying that the bubble spreads roughly ballistically. This is in agreement with the expectations from QFT, where the edge of a true vacuum is a solitonic solution expected to propagate ballistically. However, a precise quantitative statement is admittedly difficult to make. The data is reasonably well converged in bond dimension (see inset, where we plot data obtained with different bond dimensions). Yet, the bubble, a coherent collective excitation, generates long-range entanglement, making it hard to accurately capture with tensor network methods. This can be seen from the appearance of the horizontal stripes in Fig. 3(b) at later times, which are numerical artifacts. Additionally, because of the finite lattice spacing, approximating the bubble as spherical might not be very accurate at initial times. For these reasons, we view the growth of the coherent, highly-entangled bubble as a promising target for experimental quantum simulation beyond the reach of classical techniques.
@@ IV Discussion
Next, we simulated the scattering of true vacuum bubbles in a metastable false vacuum. We discovered a sharp dynamical transition between a conventional scattering regime and one where the collision induces a self-sustaining expansion of the true vacuum bubble. Calculations from quantum field theory show that this process should be exponentially suppressed. We indeed find that the threshold for inducing the false vacuum decay at small $g/J$ considerably exceeds the energy required to cross the potential barrier classically. This implies that the true vacuum bubble is hard to generate in the perturbative regime, but occurs as the non-perturbative coherent processes become available at larger $g/J$. The long-time dynamics of the bubble expansion is hard to capture classically, making this an interesting setting for quantum simulation beyond the reach of classical methods. Probing the expansion of the bubble would have implications for understanding the false vacuum decay beyond the perturbative regime but, more broadly, is a probe of collective coherent phenomena in correlated systems and thermalization dynamics of quantum metastable states [51].
From the methodological perspective, the presented work demonstrates the ability of tree tensor networks to simulate scattering dynamics on sizeable 2D lattices. This success is related to the fact that the entanglement generated in the scattering process is reasonably structured in a way that can be captured by a TTN. This should hold true in a typical scattering setup, as long as the underlying vacuum is reasonably weakly entangled. This leads us to believe that equivalent simulations could be performed in more complex models, opening the door for general simulations of scattering in two spatial dimensions.

### v4

@@ Abstract
Our results demonstrate that tensor networks can capture scattering dynamics and probe non-perturbative physics in 2D. This opens the door to such simulations in more complex settings, like lattice-gauge theories and condensed matter systems, and paves the way for future quantum simulations in regimes beyond the reach of classical simulations.
@@ I The model and its low-energy spectrum
The elementary excitations above the polarized ground state are magnons; spin flips generated by the $X$ operator. Their excitation energy is $8J$ for $g=h=0$, and they acquire a finite dispersion at finite $g/J$. At leading order the dispersion relation reads [40]:
$\varepsilon_{1}=8J-\frac{g^{2}}{J}\left(\frac{1}{2}+\frac{1}{4}\left(\cos k_{x}+\cos k_{y}\right)\right)+\mathcal{O}(g^{4})\ .$ (2)
@@ II From elastic to inelastic scattering
Figure 2: Scattering of magnon wave packets. (a) The evolution of energy density for varying $g/J$. The wave packets have momenta $\mathbf{k}=\pm\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (b) The evolution of energy density at $g/J=1.5$ for two values of $\mathbf{k}=\pm(k,k)$. (c) The evolution of the quasiparticle correlators for representative values of $g/J$. The value at $tJ=10$ is subtracted to disentangle the scattering from the background effects. (d) The $g$-dependence of the peak value of the correlators. (e) The long-range correlations. The points where the correlations are measured are sketched in the top right panel of (a) using the same colors scheme. As they are smaller than the two-point measurements, the three-point correlators are multiplied by a factor of 3 for visibility. (f) The $g$-dependence of the peak value of the long-range correlations. The curves drawn with different shades of the same color correspond to simulations with different bond dimension, the darkest being the largest. For $g/J=1.0$ we use $\chi=100$ and $150$, for $g/J=1.5$ $\chi=100,150,200$, and for $g/J>1.5$ we have $\chi=100,150,200,250$.
@@
To visualize the wave packets, we use the energy density
$\varepsilon_{i}\coloneqq-\frac{1}{2}J\sum_{j\mathrm{\ n.n.\ of\ }i}\langle Z_{i}Z_{j}\rangle-g\langle X_{i}\rangle-h\langle Z_{i}\rangle$ (3)
with $j$ nearest neighbors of $i$. Its evolution through the scattering process is shown in Fig. 2(a) for three representative values of $g/J$ and $\mathbf{k}=\pm\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. The scattering process is predominantly elastic for $g/J=1$ (top). At $g/J=1.5$ (middle) it goes through an intermediate resonance, and we find signatures of production of a heavier particle for $g/J=2.0$ (bottom).
@@
To quantify the particles produced during the scattering process, we measure correlators corresponding to approximate creation operators of composite excitations, built from products of the number operator $n_{i,j}=\frac{1}{2}\left(Z_{i,j}+\mathds{1}_{i,j}\right)$ on neighboring lattice sites. We assume that the scattering products are constrained to the resonant subspace of particles with 8 domain walls, and measure the three rightmost excitations shown in Fig. 1.
The three-spin horizontal ($h$) and kink ($k$) excitations are measured by
$\begin{split}C_{3}^{h}&=2\sum_{ij}\langle n_{i-1,j}n_{i,j}n_{i+1,j}\rangle,\\ C_{3}^{k}&=4\sum_{ij}\langle n_{i,j}n_{i+1,j}n_{i,j+1}(\mathds{1}-n)_{i+1,j+1}\rangle.\end{split}$ (4)
@@
The integer prefactors of $C_{3}^{h}$ and $C_{3}^{k}$ account for the degeneracy of these excitations; the signal of $C_{3}^{h}$ matches with the one of vertically oriented three-spin excitations, and the same for the four differently oriented three-spin kinks. The $(\mathds{1}-n)$ factor is necessary to disentangle the signals from larger particles. Without it, $C_{3}^{k}$ would also partially capture the four-spin particles described by $C_{4}$.
At finite $g$ these correlators do not measure the quasiparticles exactly, but only their approximations. One way to amend this would be to adiabatically turn off the transverse field after the scattering [10], to reinstate the product-state nature of the excitations. In contrast with 1D cases, this approach does not seem feasible in our case. The timescale for an accurate reverse quench is much larger than the decay time of heavier products, and more importantly, the time needed for the products to reach the edge of the system. Hence, we disentangle the scattering signal from the background by subtracting the value of the correlators at the end of the preparation procedure, at $tJ=10$. The raw data without the subtraction is shown in Sec. IV.5.
@@
The evolution of the correlations is shown in Fig. 2(e), and $g$-dependence of peak amplitudes in Fig. 2(f). At $g/J=1$, the single dominant peak is a signature of the elastic process. It splits into two with increasing $g$, and at $g/J\gtrsim 1.5$ we clearly find two separate contributions. The first still corresponds to the elastic process, while the second signals the delayed arrival of particles generated through an intermediate resonance. The second peak is accompanied by a peak in $c_{O}$. The peaks of $c_{A}$ and $c_{O}$ are not equal because the magnon dispersion is not exactly sinusoidal, and thus the scattering is not precisely at the van Hove singularity. Note that the products are more concentrated in the off-axis direction for $g/J=1.5$. We expect that this is accompanied by a finite-size effect, as the products are not point-like particles, but rather extended objects on the lattice. The negative $c_{AO}$ that accompanies the overlap of $c_{A}$ and $c_{O}$ shows that the presence of particles on the two axes is anti-correlated. This confirms that these are two separate two-particle processes, and not a four-particle process.
@@ III Scattering in a false vacuum
Figure 3: Scattering in the false vacuum. (a) A sketch of the energy landscape of a true vacuum bubble with radius $r$. (b) The evolution of the energy density during scattering with $h/J=0.1$ and $h/J=0.3$, both with $g/J=1.5$ and $\mathbf{k}=\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (c) The change in magnetization per site with varying $h$ for $g/J=1.5$ and $\mathbf{k}=\left(\frac{\pi}{2},\frac{\pi}{2}\right)$. (d) The dependence of the threshold value $h^{*}$ on $g/J$. The gray dashed line corresponds to the threshold $h^{*}$ expected for $g\rightarrow 0$. See text for details. (e) The evolution of the radius of the true vacuum bubble in log-log scale for varying $g$, with some longitudinal field larger than the threshold value. We subtract the time $t_{0}$ when the radius reaches $r_{0}=3$ on the horizontal, and $r_{0}$ on the vertical axis. The black dashed line corresponds to linear growth and acts as a guide to the eye. (Inset) The evolution of the radius in lin-lin scale. In (c) and (e), the curves with different shades of the same color correspond to simulations with different bond dimension, darker is bigger. Here we use $\chi=150$ and 200.
@@
Originating in cosmology and high-energy physics [42], the question has recently attracted attention in the context of metastability in quantum many-body systems. It has been studied in 1D numerically [43, 44, 28, 45, 46, 47, 48] and experimentally [49, 50, 51, 52]. The microscopic mechanism behind false vacuum decay is probabilistic bubble nucleation [45]: the vacuum fluctuations induced through a quench are small bubbles of the true vacuum, where the system locally tunnels to the stable state. Assuming a spherical bubble with radius $r$, its energy consists of two contributions: a positive interface tension, and a negative bulk term coming from the energy difference between the two vacua [42, 45]. In 2D one expects the energy $E(r)=A\cdot 2J\cdot r-B\cdot 2h\cdot r^{2}$, with model-dependent geometrical factors $A$ and $B$. The parabola acts as a potential barrier between the false vacuum at $r=0$ and the true vacuum at $r\rightarrow\infty$, with the maximal energy at the critical bubble size $r^{*}\propto J/h$ and $E(r^{*})\propto J^{2}/h$. If a bubble bigger than $r^{*}$ is created, it can expand by transforming the potential energy of the false vacuum into the kinetic energy of the bubble’s walls. The system thus transitions towards the stable manifold. The energy manifold is sketched in Fig. 3(a).
Here, we prepare magnon wave packets in a false vacuum background, and study how their collision induces the transition into the stable manifold. In this case, we do not expect the system to fully tunnel through the potential barrier as in unperturbed false vacuum decay, but rather over it. Intuitively, the collision of the wave packets concentrates energy in the scattering site, and if the energy density is large enough, the system locally traverses the potential barrier and creates an expanding critical bubble.
@@
Because the height of the potential barrier $E(r^{*})\propto 1/h$, we can vary it by varying $h$ at fixed $g$. By doing this, we find two sharply separated dynamical regimes; see the evolution of energy density in Fig. 3(b) for two representative values of $h$. At the smaller $h$ we recover the process akin to the one in Fig. 2. However, if $h$ is tuned beyond a threshold value, denoted $h^{*}$, the dynamics drastically change. Here, the scattering process generates a bubble of the true vacuum, which violently expands and encompasses the whole system. This is reflected in the evolution of the magnetization, shown in Fig. 3(c), exhibiting two types of behavior depending on $h$: it is approximately constant at $h<h^{*}$, and rapidly increases after scattering when $h>h^{*}$.
In Fig. 3(d), we show the dependence of the threshold $h^{*}$ on $g$. We obtain the points by simulating the scattering process with a set $g$ and varying $h$, and denoting $h^{*}$ as the smallest value of $h$ where we observe a large change of average magnetization. The solid line is a guide to the eye. In the limit of $g\rightarrow 0$, where the model is classical, we can exactly determine the energies involved. By equating the energy of the wave packets $2\times 8J$ with $E(r=r^{*},h^{*})$ (assuming a spherical bubble in 2D; $A=2\pi$, $B=\pi$), we find $h^{*}(g\rightarrow 0)\sim 0.39$, denoted by the gray dashed line in Fig. 3(d). Interestingly, this value strongly disagrees with the simulations. This indicates that the creation of the bubble is not conditioned only on the energetics of the incoming particles, but also on the nature of their interaction, which depends on $g/J$. In particular, the critical bubble is a many-body excitation, and $g/J$ needs to be large enough to enable its non-perturbative generation from the initial two-particle state.
Induced false vacuum decay has been studied in the context of quantum field theory in Refs. [36, 37, 38]. These works study the process in the semi-classical limit, where it is shown that a collision of two particles does not trigger the decay: the scattering cross-section for such process remains exponentially suppressed up to infinite energy. A simple explanation is that while the scattering might carry enough energy to create the true vacuum bubble, it is a state of many excitations. Generating it from a two-particle state is an unlikely high-order process. This is in accordance with our results, which show that the process is suppressed at small $g/J$ even when enough energy is available. We however find that increasing $g/J$ beyond the perturbative regime allows us to induce the critical bubble.
On the other hand, the concentration of the energy in the collision point locally heats the system, and the decay process might be related to false vacuum decay at finite-temperature [53]. This is a classical process in which thermal fluctuations bring the system over the potential barrier. However, this would mean only a weak dependence on $g/J$ (only due to the renormalization of the domain walls), whilst we observe a strong effect in Fig. 3(d).
Additionally, the presented results are on a finite lattice, and frankly quite far from the continuum limit of field theory. While scaling these simulations to much larger system sizes is a formidable task, we rule out the lattice effects by checking that the threshold $h^{*}$ does not depend on the size of the wave packet. See Sec. IV.6.
@@
Finally, in Fig. 3(e), we plot the evolution of the bubble radius for various $g/J$, with $h>h^{*}(g)$. The main panel shows a log-log plot where we subtract a $g$-dependent $t_{0}$ in the time axis, and $r_{0}=3$ in the vertical axis. The inset shows the raw data in linear scale. The black dashed line corresponds to linear increase of $r$. It roughly matches the numerical data, implying that the bubble spreads ballistically, in agreement to a solitonic solution expected in QFT. However, a precise quantitative statement is admittedly difficult to make. The data is reasonably well converged in bond dimension (see inset, where we plot data obtained with different bond dimensions). Yet, the bubble, a coherent collective excitation, generates long-range entanglement, making it hard to accurately capture with tensor network methods. This can be seen from the appearance of the horizontal stripes in Fig. 3(b) at later times, which are numerical artifacts. Additionally, because of the finite lattice spacing, approximating the bubble as spherical might not be very accurate at initial times. For these reasons, we view the growth of the coherent, highly-entangled bubble as a promising target for experimental quantum simulation beyond the reach of classical techniques.
In conclusion, we believe that the observed phenomenon is due to a combination of a local thermal and a quantum tunneling process. Further work is certainly required to completely understand their interplay. It should be noted that it is not known whether Coleman’s arguments and the critical droplet theory even apply in two dimensions [54]. We plan to investigate this in an upcoming work.
@@ IV Discussion
Next, we simulated the scattering of true vacuum bubbles in a metastable false vacuum. We discovered a sharp dynamical transition between a conventional scattering regime and one where the collision induces a self-sustaining expansion of the true vacuum bubble. We find that the threshold for inducing the false vacuum decay at small $g/J$ considerably exceeds the energy required to cross the potential barrier classically. This leads us to believe that the generation of the critical bubble is a collective non-perturbative process. The long-time dynamics of the bubble expansion is hard to capture classically, making this an interesting setting for quantum simulation beyond the reach of classical methods. Probing the expansion of the bubble would have implications for understanding the false vacuum decay beyond the perturbative regime but, more broadly, is a probe of collective coherent phenomena in correlated systems and thermalization dynamics of quantum metastable states [55].
From the methodological perspective, the presented work demonstrates that tree tensor networks can simulate scattering dynamics on sizeable 2D lattices. This success is related to the fact that the entanglement generated in the scattering process is reasonably structured in a way that can be captured by a TTN. This should hold true in a typical scattering setup, as long as the underlying vacuum is reasonably weakly entangled. This leads us to believe that equivalent simulations could be performed in more complex models, opening the door for general simulations of scattering in two spatial dimensions.